percentage by mole of NO2 in a mixture of NOV and NO having average molecular mass 34 is

Grade:11

1 Answers

Vikas TU
7 years ago
Let x and y be the mole fraction of NO2 and NO in the mixture.
we know x+y = 1
and from molecular weight formulae:
M1*x + M2*y = (40+36)/2 = 34.
we know M1 = 46
and M2 = 30.
Solve the eqns. and get x and y values.